# Towards Aligning Language Models with Textual Feedback

### v1

@@ 6 Conclusion
We presented ALT, an approach that uses textual feedback to align an LM. Our findings across diverse tasks, such as reducing model toxicity, improving summarization, and aligning dialogue, underscore the efficacy and efficiency of this approach. Notably, ALT surpasses traditional reinforcement learning methods like PPO in toxicity reduction and achieves comparable summarization performance with considerably fewer training samples. Furthermore, our results indicate the feasibility of leveraging large language models to provide effective feedback for aligning dialogue models. Our current experiments failed to show improvements with more detailed textual feedback. However, we believe that this outcome could change with more consistent feedback. We leave this exploration for future work. Our findings open promising directions for further research into using varied types of feedback to improve LLM alignment.
@@ A.2 Summarization
We fine-tune the GPT-J SFT model using the language feedback mentioned on 3.4.2. As the distinct phases of our algorithm are decoupled, one can leverage different computation resources at every stage. The data collection stage is the most costly one in terms of time required to sample and provide feedback to several generations, but one can launch multiple smaller GPU jobs and employ frameworks for faster inference such as vLLM Kwon et al. (2023) for substantial speedup. We carried out the sampling phase on 8 parallel NVIDIA 3090 (24GB) threads that sampled on different disjoint subsets of the 2048 prompts. Following this setup, the $2048*96=196608$ generations can be sampled in less than 10min. The same can be applied for the feedback stage, either for Quark and $\textsc{ALT}_{\textsc{RM}}$, which only require running the Reward Model and can be done in several independent NVIDIA 3090 GPUs; or for $\textsc{ALT}_{\textsc{LM}}$ in which several independent CPU-only jobs can query the GPT-3.5 model through the OpenAI API. For all experiments, training is done in 2 NVIDIA A100 (80GB) and employing DeepSpeed Rasley et al. (2020) for handling the training parallelization and leveraging the ZeRO-optimizer and CPU offloading features. Each training iteration takes  3-4h to complete.
@@ A.3 Dialogue
We fine-tune the Pythia-2.8b SFT model using the language feedback mentioned on 3.4.3. The same decoupled nature for the data collection and training described in A.2 applies here, and we leveraged the same computation resources. Figure 5 contains the generations’ length and % of truncated generations along iterations for $\textsc{ALT}_{\textsc{LMC}}$ and SteerLM, and Figure 6 contains the perplexity curve as training progresses.

### v2

@@ 6 Conclusion
We presented ALT, an approach that uses textual feedback to align an LLM. Our findings across diverse tasks, such as reducing model toxicity, improving summarization, and aligning dialogue, underscore the efficacy and efficiency of this approach. Notably, ALT surpasses traditional reinforcement learning methods like PPO in toxicity reduction and achieves comparable summarization performance with considerably fewer training samples. Furthermore, our results indicate the feasibility of leveraging large language models to provide effective feedback for aligning dialogue models. Our current experiments failed to show improvements with more detailed textual feedback. However, we believe that this outcome could change with more consistent feedback. Our findings open promising directions for further research into the use of varied types of feedback to improve LLM alignment.
@@
Acknowledgments
This project was funded by a grant from the Swiss National Science Foundation (Project No. 197155), a Responsible AI grant by the Haslerstiftung and an IBM PhD fellowship to SD.
@@ A.2 Summarization
We fine-tune the GPT-J SFT model using the language feedback mentioned in subsection 3.2. As the distinct phases of our algorithm are decoupled, one can leverage different computation resources at every stage. The data collection stage is the most costly one in terms of time required to sample and provide feedback to several generations, but one can launch multiple smaller GPU jobs and employ frameworks for faster inference such as vLLM Kwon et al. (2023) for substantial speedup. We carried out the sampling phase on 8 parallel NVIDIA 3090 (24GB) threads that sampled on different disjoint subsets of the 2048 prompts. Following this setup, the $2048*96=196608$ generations can be sampled in less than 10min. The same can be applied for the feedback stage, either for Quark and $\textsc{ALT}_{\textsc{RM}}$, which only require running the Reward Model and can be done in several independent NVIDIA 3090 GPUs; or for $\textsc{ALT}_{\textsc{LM}}$ in which several independent CPU-only jobs can query the GPT-3.5 model through the OpenAI API. For all experiments, training is done in 2 NVIDIA A100 (80GB) and employing DeepSpeed Rasley et al. (2020) for handling the training parallelization and leveraging the ZeRO-optimizer and CPU offloading features. Each training iteration takes  3-4h to complete.
@@ A.3 Dialogue
We fine-tune the Pythia-2.8b SFT model using the language feedback mentioned in subsection 3.3. The same decoupled nature for the data collection and training described in A.2 applies here, and we leveraged the same computation resources. Figure 5 contains the generations’ length and % of truncated generations along iterations for $\textsc{ALT}_{\textsc{LMC}}$ and SteerLM, and Figure 6 contains the perplexity curve as training progresses.